# Tree Search for Language Model Agents

### v2

@@ A.3.1 Language Model Agents
example_user: IMAGES: (1) current page screenshot OBSERVATION: [31] [IMG] [Image, description: hp fx-7010dn fax machine, url: B08GKZ3ZKD.0.jpg] [32] [A] [HP CB782A#ABA 640 Inkjet Fax Machine (Renewed)] [ ] [StaticText] [\$279.49] [33] [BUTTON] [Add to Cart] [34] [A] [Add to Wish List] [35] [A] [Add to Compare] URL: http://onestopmarket.com/office-products/office-electronics.html OBJECTIVE: What is the price of HP Inkjet Fax Machine? PREVIOUS ACTION: None example_assistant: Let’s think step-by-step. This page list the information of HP Inkjet Fax Machine, which is the product identified in the objective. Its price is \$279.49. I think I have achieved the objective. I will issue the stop action with the answer. In summary, the next action I will perform is ´´´stop [\$279.49]´´´ example_user: IMAGES: (1) current page screenshot OBSERVATION: [ ] [StaticText] [/f/food] [ ] [StaticText] [[homemade] Obligatory Halloween Pumpkin Loaf!   Submitted by  kneechalice  t3_yid9lu  1 year ago] [9] [IMG] [ ] [ ] [StaticText] [Submitted by  kneechalice  t3_yid9lu  1 year ago] [10] [A] [kneechalice] [11] [A] [45 comments] [ ] [StaticText] [[I ate] Maple Pecan Croissant   Submitted by  AccordingtoJP  t3_y3hrpn  1 year ago] [14] [IMG] [ ] [ ] [StaticText] [Submitted by  AccordingtoJP  t3_y3hrpn  1 year ago] [15] [A] [AccordingtoJP] [16] [A] [204 comments] URL: http://reddit.com OBJECTIVE: Tell me what the top comment on the croissant post says. PREVIOUS ACTION: None example_assistant: Let’s think step-by-step. This page has a post titled ’[I ate] Maple Pecan Croissant’, which is the post mentioned in the objective. In order to find the top comment, I will navigate into the comments section of the post. In summary, the next action I will perform is ´´´click [11]´´´ example_user: IMAGES: (1) current page screenshot OBSERVATION: [ ] [StaticText] [What are you looking for today?] [5] [INPUT] [ ] [6] [SELECT] [Select a category] [7] [BUTTON] [Search] [ ] [StaticText] [Latest Listings] [ ] [StaticText] [Atlas Powered Audio System w/ Tripod  150.00 \$  Music instruments  Borough of Red Lion  (Pennsylvania)  2023/11/16] [8] [IMG] [Atlas Powered Audio System w/ Tripod] [9] [A] [Atlas Powered Audio System w/ Tripod] [ ] [StaticText] [150.00 \$] [ ] [StaticText] [Neptune Gaming Console  350.00 \$  Video gaming  Pennwyn  (Pennsylvania)  2023/11/16] [10] [IMG] [Neptune Gaming Console] [11] [A] [Neptune Gaming Console] [ ] [StaticText] [350.00 \$] URL: http://classifieds.com OBJECTIVE: Help me find the cheapest dark colored guitar. PREVIOUS ACTION: None example_assistant: Let’s think step-by-step. The objective is to find the cheapest dark colored guitar on the site. The site has a search box whose ID is [5]. I can search for guitars by entering “guitar”. I can submit this by pressing the Enter afterwards. In summary, the next action I will perform is ´´´type [5] [guitar]  [1]´´´
@@
example_user: OBSERVATION: [1744] [link] [’HP CB782A#ABA 640 Inkjet Fax Machine (Renewed)’] [1749] [StaticText] [\$279.49] [1757] [button] [’Add to Cart’] [1760] [button] [’Add to Wish List’] [1761] [button] [’Add to Compare’] URL: http://onestopmarket.com/office-products/office-electronics.html OBJECTIVE: What is the price of HP Inkjet Fax Machine? PREVIOUS ACTION: None example_assistant: Let’s think step-by-step. This page list the information of HP Inkjet Fax Machine, which is the product identified in the objective. Its price is \$279.49. I think I have achieved the objective. I will issue the stop action with the answer. In summary, the next action I will perform is ´´´stop [\$279.49]´´´ example_user: IMAGES: (1) current page screenshot OBSERVATION: [204] [heading] [’/f/food’] [593] [heading] [’[homemade] Obligatory Halloween Pumpkin Loaf!’] [942] [link] [’[homemade] Obligatory Halloween Pumpkin Loaf!’] [945] [StaticText] [’Submitted by ’] [30] [link] [’kneechalice’ expanded: False] [1484] [StaticText] [’t3_yid9lu’] [949] [time] [’October 31, 2022 at 10:10:03 AM EDT’] [1488] [StaticText] [’1 year ago’] [1489] [link] [’45 comments’] [605] [heading] [’[I ate] Maple Pecan Croissant’] [963] [link] [’[I ate] Maple Pecan Croissant’] [966] [StaticText] [’Submitted by ’] [37] [link] [’AccordingtoJP’ expanded: False] [1494] [StaticText] [’t3_y3hrpn’] [970] [time] [’October 13, 2022 at 10:41:09 PM EDT’] [1498] [StaticText] [’1 year ago’] [1499] [link] [’204 comments’] URL: http://reddit.com
@@
PREVIOUS ACTION: None example_assistant: Let’s think step-by-step. This page has a post titled ’[I ate] Maple Pecan Croissant’, which is the post mentioned in the objective. In order to find the top comment, I will navigate into the comments section of the post. In summary, the next action I will perform is ´´´click [1499]´´´ example_user: IMAGES: (1) current page screenshot OBSERVATION: [42] [link] [’My account’] [43] [link] [’Logout’] [44] [link] [’Publish Ad’] [25] [heading] [’What are you looking for today?’] [143] [StaticText] [’Keyword’] [81] [textbox] [’e.g., a blue used car’ required: False] [146] [StaticText] [’Category’] [28] [heading] [’Latest Listings’] [86] [link] [’Atlas Powered Audio System w/ Tripod’] [176] [img] [’Atlas Powered Audio System w/ Tripod’] [511] [StaticText] [’150.00 \$’] [88] [link] [’Neptune Gaming Console’] [178] [img] [’Neptune Gaming Console’] [515] [StaticText] [’350.00 \$’] URL: http://classifieds.com
@@
PREVIOUS ACTION: None example_assistant: Let’s think step-by-step. The objective is to find the cheapest dark colored guitar on the site. The site has a search box whose ID is [5]. I can search for guitars by entering “guitar”. I can submit this by pressing the Enter afterwards. In summary, the next action I will perform is ´´´type [81] [guitar]  [1]´´´
@@
example_user: OBSERVATION: [1744] [link] [’HP CB782A#ABA 640 Inkjet Fax Machine (Renewed)’] [1749] [StaticText] [\$279.49] [1757] [button] [’Add to Cart’] [1760] [button] [’Add to Wish List’] [1761] [button] [’Add to Compare’] URL: http://onestopmarket.com/office-products/office-electronics.html OBJECTIVE: What is the price of HP Inkjet Fax Machine? PREVIOUS ACTION: None example_assistant: Let’s think step-by-step. This page list the information of HP Inkjet Fax Machine, which is the product identified in the objective. Its price is \$279.49. I think I have achieved the objective. I will issue the stop action with the answer. In summary, the next action I will perform is ´´´stop [\$279.49]´´´ example_user: IMAGES: (1) current page screenshot OBSERVATION: [164] [textbox] [’Search’ focused: True required: False] [171] [button] [’Go’] [174] [link] [’Find directions between two points’] [212] [heading] [’Search Results’] [216] [button] [’Close’] URL: http://openstreetmap.org
@@
PREVIOUS ACTION: None example_assistant: Let’s think step-by-step. This page has a search box whose ID is [164]. According to the nominatim rule of openstreetmap, I can search for the restaurants near a location by “restaurants near”. I can submit my typing by pressing the Enter afterwards. In summary, the next action I will perform is ´´´ type [164][restaurants near CMU][1]´´´
@@ A.3.2 Value Function
There are three types of tasks: 1. Information seeking: The user wants to obtain certain information from the webpage, such as the information of a product, reviews, the text in a comment or post, the date of a submission, etc. This may be formulated in the intent as “tell me”, “what is”, or “list out”. The agent’s response must contain the information the user wants, or explicitly state that the information is not available. Otherwise, e.g. the agent encounters an exception and respond with the error content, the task is considered to be a failure. It is VERY IMPORTANT that the bot response is the stop action with the correct output. If the bot response is not stop (e.g., it is click, type, or goto), it is considered a failure for information seeking tasks. 2. Site navigation: The user wants to navigate to a specific page (which may also be specified in the intent as “find”, “show me”, “navigate to”). Carefully examine the agent’s action history and the final state of the webpage (shown in the LAST IMAGE) to determine whether the agent successfully completes the task. It is VERY IMPORTANT that the agent actually navigates to the specified page (reflected by the final state of the webpage, in the LAST IMAGE) and NOT just output the name of the item or post. Make sure that the final url is compatible with the task. For example, if you are tasked to navigate to a comment or an item, the final page and url should be that of the specific comment/item and not the overall post or search page. If asked to navigate to a page with a similar image, make sure that an image on the page is semantically SIMILAR to the intent image. If asked to look for a particular post or item, make sure that the image on the page is EXACTLY the intent image. For this type of task to be considered successful, the LAST IMAGE and current URL should reflect the correct content. No need to consider the agent’s response. 3. Content modification: The user wants to modify the content of a webpage or configuration. Ensure that the agent actually commits to the modification. For example, if the agent writes a review or a comment but does not click post, the task is considered to be a failure. Carefully examine the agent’s action history and the final state of the webpage to determine whether the agent successfully completes the task. No need to consider the agent’s response. *IMPORTANT* Format your response into two lines as shown below: Thoughts: <your thoughts and reasoning process> Status: “success” or “failure” On the right track to success: “yes” or “no” user:
@@
Bot response to the user: last_response Current URL: current_url The images corresponding to the user intent are shown in the FIRST {len(intent_images)} images (before the User Intent). The last {len(screenshots)} snapshots of the agent’s trajectory are shown in the LAST {len(screenshots)} images. The LAST IMAGE represents the current state of the webpage.
@@
  • Rather than a binary output, we instruct the model to produce whether the given observations have succeeded at the task or failed. If it fails, we further prompt the model to output if it is possibly on the right track to success. This allows us to collect scores in `$\{0,0.5,1\}$, enabling more finegrained value outputs (in addition to the averaging of multiple reasoning paths described in Sec. 4.1).
@@ A.4 Search Algorithm
Our search procedure described in Sec. 3.3 is summarized in Algorithm 1.

### v3

@@ 1. You should only issue an action that is valid given the current observation
example_user:
IMAGES: (1) current page screenshot
OBSERVATION:
[31] [IMG] [Image, description: hp fx-7010dn fax machine, url: B08GKZ3ZKD.0.jpg]
[32] [A] [HP CB782A#ABA 640 Inkjet Fax Machine (Renewed)]
[ ] [StaticText] [\$279.49]
[33] [BUTTON] [Add to Cart]
[34] [A] [Add to Wish List]
[35] [A] [Add to Compare]
URL: http://onestopmarket.com/office-products/office-electronics.html
OBJECTIVE: What is the price of HP Inkjet Fax Machine?
PREVIOUS ACTION: None
example_assistant: Let’s think step-by-step. This page list the information of HP Inkjet Fax Machine, which is the product identified in the objective. Its price is \$279.49. I think I have achieved the objective. I will issue the stop action with the answer. In summary, the next action I will perform is ´´´stop [\$279.49]´´´
example_user:
IMAGES: (1) current page screenshot
OBSERVATION:
[ ] [StaticText] [/f/food]
[ ] [StaticText] [[homemade] Obligatory Halloween Pumpkin Loaf! Submitted by kneechalice t3_yid9lu 1 year ago]
[9] [IMG] [ ]
[ ] [StaticText] [Submitted by kneechalice t3_yid9lu 1 year ago]
[10] [A] [kneechalice]
[11] [A] [45 comments]
[ ] [StaticText] [[I ate] Maple Pecan Croissant Submitted by AccordingtoJP t3_y3hrpn 1 year ago]
[14] [IMG] [ ]
[ ] [StaticText] [Submitted by AccordingtoJP t3_y3hrpn 1 year ago]
[15] [A] [AccordingtoJP]
[16] [A] [204 comments]
URL: http://reddit.com
OBJECTIVE: Tell me what the top comment on the croissant post says.
PREVIOUS ACTION: None
example_assistant: Let’s think step-by-step. This page has a post titled ’[I ate] Maple Pecan Croissant’, which is the post mentioned in the objective. In order to find the top comment, I will navigate into the comments section of the post. In summary, the next action I will perform is ´´´click [11]´´´
example_user:
IMAGES: (1) current page screenshot
OBSERVATION:
[ ] [StaticText] [What are you looking for today?]
[5] [INPUT] [ ]
[6] [SELECT] [Select a category]
[7] [BUTTON] [Search]
[ ] [StaticText] [Latest Listings]
[ ] [StaticText] [Atlas Powered Audio System w/ Tripod 150.00 \$ Music instruments Borough of Red Lion (Pennsylvania) 2023/11/16]
[8] [IMG] [Atlas Powered Audio System w/ Tripod]
[9] [A] [Atlas Powered Audio System w/ Tripod]
[ ] [StaticText] [150.00 \$]
[ ] [StaticText] [Neptune Gaming Console 350.00 \$ Video gaming Pennwyn (Pennsylvania) 2023/11/16]
URL: http://classifieds.com
OBJECTIVE: Help me find the cheapest dark colored guitar.
PREVIOUS ACTION: None
example_assistant: Let’s think step-by-step. The objective is to find the cheapest dark colored guitar on the site. The site has a search box whose ID is [5]. I can search for guitars by entering “guitar”. I can submit this by pressing the Enter afterwards. In summary, the next action I will perform is ´´´type [5] [guitar] [1]´´´
@@
example_user:
OBSERVATION:
[1744] [link] [’HP CB782A#ABA 640 Inkjet Fax Machine (Renewed)’]
[1749] [StaticText] [\$279.49]
[1757] [button] [’Add to Cart’]
[1760] [button] [’Add to Wish List’]
[1761] [button] [’Add to Compare’]
URL: http://onestopmarket.com/office-products/office-electronics.html
OBJECTIVE: What is the price of HP Inkjet Fax Machine?
PREVIOUS ACTION: None
example_assistant: Let’s think step-by-step. This page list the information of HP Inkjet Fax Machine, which is the product identified in the objective. Its price is \$279.49. I think I have achieved the objective. I will issue the stop action with the answer. In summary, the next action I will perform is ´´´stop [\$279.49]´´´ example_user:
IMAGES: (1) current page screenshot
OBSERVATION:
[204] [heading] [’/f/food’]
[593] [heading] [’[homemade] Obligatory Halloween Pumpkin Loaf!’]
[942] [link] [’[homemade] Obligatory Halloween Pumpkin Loaf!’]
[945] [StaticText] [’Submitted by ’]
[30] [link] [’kneechalice’ expanded: False]
[1484] [StaticText] [’t3_yid9lu’]
[949] [time] [’October 31, 2022 at 10:10:03 AM EDT’]
[1488] [StaticText] [’1 year ago’]
[1489] [link] [’45 comments’]
[605] [heading] [’[I ate] Maple Pecan Croissant’]
[963] [link] [’[I ate] Maple Pecan Croissant’]
[966] [StaticText] [’Submitted by ’]
[37] [link] [’AccordingtoJP’ expanded: False]
[1494] [StaticText] [’t3_y3hrpn’]
[970] [time] [’October 13, 2022 at 10:41:09 PM EDT’]
[1498] [StaticText] [’1 year ago’]
[1499] [link] [’204 comments’]
URL: http://reddit.com
@@
PREVIOUS ACTION: None
example_assistant: Let’s think step-by-step. This page has a post titled ’[I ate] Maple Pecan Croissant’, which is the post mentioned in the objective. In order to find the top comment, I will navigate into the comments section of the post. In summary, the next action I will perform is ´´´click [1499]´´´ example_user:
IMAGES: (1) current page screenshot
OBSERVATION:
[42] [link] [’My account’]
[43] [link] [’Logout’]
[44] [link] [’Publish Ad’]
[25] [heading] [’What are you looking for today?’]
[143] [StaticText] [’Keyword’]
[81] [textbox] [’e.g., a blue used car’ required: False]
[146] [StaticText] [’Category’]
[28] [heading] [’Latest Listings’]
URL: http://classifieds.com
@@
PREVIOUS ACTION: None
example_assistant: Let’s think step-by-step. The objective is to find the cheapest dark colored guitar on the site. The site has a search box whose ID is [5]. I can search for guitars by entering “guitar”. I can submit this by pressing the Enter afterwards. In summary, the next action I will perform is ´´´type [81] [guitar] [1]´´´
@@
example_user:
OBSERVATION:
[1744] [link] [’HP CB782A#ABA 640 Inkjet Fax Machine (Renewed)’]
[1749] [StaticText] [\$279.49]
[1757] [button] [’Add to Cart’]
[1760] [button] [’Add to Wish List’]
[1761] [button] [’Add to Compare’]
URL: http://onestopmarket.com/office-products/office-electronics.html
OBJECTIVE: What is the price of HP Inkjet Fax Machine?
PREVIOUS ACTION: None
example_assistant: Let’s think step-by-step. This page list the information of HP Inkjet Fax Machine, which is the product identified in the objective. Its price is \$279.49. I think I have achieved the objective. I will issue the stop action with the answer. In summary, the next action I will perform is ´´´stop [\$279.49]´´´ example_user:
IMAGES: (1) current page screenshot
OBSERVATION:
[164] [textbox] [’Search’ focused: True required: False]
[171] [button] [’Go’]
[174] [link] [’Find directions between two points’]
[212] [heading] [’Search Results’]
[216] [button] [’Close’]
URL: http://openstreetmap.org
@@
PREVIOUS ACTION: None
example_assistant: Let’s think step-by-step. This page has a search box whose ID is [164]. According to the nominatim rule of openstreetmap, I can search for the restaurants near a location by “restaurants near”. I can submit my typing by pressing the Enter afterwards. In summary, the next action I will perform is ´´´ type [164][restaurants near CMU][1]´´´
@@ A.3.2 Value Function
There are three types of tasks:
1. Information seeking: The user wants to obtain certain information from the webpage, such as the information of a product, reviews, the text in a comment or post, the date of a submission, etc. This may be formulated in the intent as “tell me”, “what is”, or “list out”. The agent’s response must contain the information the user wants, or explicitly state that the information is not available. Otherwise, e.g. the agent encounters an exception and respond with the error content, the task is considered to be a failure. It is VERY IMPORTANT that the bot response is the stop action with the correct output. If the bot response is not stop (e.g., it is click, type, or goto), it is considered a failure for information seeking tasks.
2. Site navigation: The user wants to navigate to a specific page (which may also be specified in the intent as “find”, “show me”, “navigate to”). Carefully examine the agent’s action history and the final state of the webpage (shown in the LAST IMAGE) to determine whether the agent successfully completes the task. It is VERY IMPORTANT that the agent actually navigates to the specified page (reflected by the final state of the webpage, in the LAST IMAGE) and NOT just output the name of the item or post. Make sure that the final url is compatible with the task. For example, if you are tasked to navigate to a comment or an item, the final page and url should be that of the specific comment/item and not the overall post or search page. If asked to navigate to a page with a similar image, make sure that an image on the page is semantically SIMILAR to the intent image. If asked to look for a particular post or item, make sure that the image on the page is EXACTLY the intent image. For this type of task to be considered successful, the LAST IMAGE and current URL should reflect the correct content. No need to consider the agent’s response.
3. Content modification: The user wants to modify the content of a webpage or configuration. Ensure that the agent actually commits to the modification. For example, if the agent writes a review or a comment but does not click post, the task is considered to be a failure. Carefully examine the agent’s action history and the final state of the webpage to determine whether the agent successfully completes the task. No need to consider the agent’s response.
*IMPORTANT*
Format your response into two lines as shown below:
Thoughts: <your thoughts and reasoning process>
Status: “success” or “failure”
On the right track to success: “yes” or “no”
user:
@@
Bot response to the user: last_response
Current URL: current_url
The images corresponding to the user intent are shown in the FIRST {len(intent_images)} images (before the User Intent).
The last {len(screenshots)} snapshots of the agent’s trajectory are shown in the LAST {len(screenshots)} images. The LAST IMAGE represents the current state of the webpage.
@@
  • Rather than a binary output, we instruct the model to produce whether the given observations have succeeded at the task or failed. If it fails, we further prompt the model to output if it is possibly on the right track to success. This allows us to collect scores in $\{0,0.5,1\}$, enabling more finegrained value outputs (in addition to the averaging of multiple reasoning paths described in Sec. 4.1).
@@ A.4 Search Algorithm
Our search procedure described in Sec. 3.3 is summarized in Algorithm. 1.
A.4.1 Environment Reset
In this section, we describe the implementation details of the backtracking used in our search procedure:
  1. 1.
    We maintain a max priority queue that contains sequences of actions and their score $v$ (from the value function). Each element is a sequence of actions that the agent has to sequentially execute starting from the initial state (task dependent, but often the website homepage) to get to state $s$ that has the corresponding score $v$.
  2. 2.
    After we execute a new action (L23 of Algorithm. 1), we append this action to the sequence of actions and add the new sequence to the priority queue with its corresponding score $v$.
  3. 3.
    In order to reset the environment to get a clean slate for the next node to explore, we reset to the initial state again, and repeat the execution of the next sequence of actions starting from step 1.
We implemented backtracking in this fashion, as we found that this was a substantially more complete way of resetting the state, as opposed to simply clicking the “back” button on the browser for example, as this does not persist certain web states such as the scroll offset, or retain text in text inputs. While our implementation does improve fidelity of backtracking and resets, it however does add significant overhead in terms of time (see Sec. 5.4 for more discussion).